b. How long will a call be responded if it has a z-score of – 1.83?


1
Expert's answer
2022-04-06T16:53:01-0400

z = -1.83 , μ = 30min, δ =4.5min,

x = zδ +μx = -1.83(4.5) + 30min =-8.325 +30 = 21.675min